OpenStudy (anonymous):

Jim is selling hot dogs at a ball game. It cost Jim $250 to purchase everything to make the hot dogs. Jim sells hot dogs for $2.00 each. If he sells hot dogs, write the equation that models his profit? Do not use spaces in your answer. You must use P and h for this equation as the variables.

OpenStudy (anonymous):

how many hotdoys does he sell?

OpenStudy (anonymous):

250-2x i guess x being the number of hotdogs sold

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

you guess ?

OpenStudy (anonymous):

do you have any idea on how many hotdogs where sold?

OpenStudy (anonymous):

cause basically just put that value in place of x and you get the profit

OpenStudy (anonymous):

It does not say

OpenStudy (anonymous):

but i am assuming you have to make an equation so i put x in there

OpenStudy (anonymous):

P=250-2h

OpenStudy (anonymous):

that looks right , I'll try it

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

we got a pro

OpenStudy (whpalmer4):

Profit = revenue - cost His cost is $250 to buy the hot dogs and everything he needs Revenue = number of hot dogs sold * sales price per hot dog If we let \(h\) be the number sold, revenue = \(2.00 h\) Profit will be given by \[P = 2.00h - 250\] If he doesn't sell any, then his profit is -250 — the cost of all the hot dogs and so on. If he sells 125 hot dogs, his profit is 2*125 - 250 = 0 — he breaks even If he sells more than 125 hot dogs, he turns a profit.

OpenStudy (whpalmer4):

see how that fits together?

OpenStudy (anonymous):

ohhhhh. That makes more sence! thanks for the explanation.

OpenStudy (whpalmer4):

They probably want "P=2h-250" as the form of the answer.

OpenStudy (whpalmer4):

"sense" :-)

OpenStudy (whpalmer4):

Always good after you build an equation to test it out and see if the results make sense. In this case, the earlier P=250-2h looks promising, until you realize that it starts out with a $250 profit before you sell any hot dogs, and you make less and less money as you sell more hot dogs :-)

OpenStudy (whpalmer4):

Which is not to say that unskilled business practitioners never unintentionally do that! :-)

OpenStudy (anonymous):

lol